# A Unified Dark-Matter–Driven Relativistic Bondi Route to Black-Hole Growth from Stellar to Supermassive Scales

### v1

@@ Acknowledgements.
In this supplementary material, we will first discuss the possibility of forming the super-Eddington supermassive black holes (SMBH), which we mean the black holes with masses $\geq 10^{9}M_{\odot}$ formed around the redshift $z\simeq 6\textrm{--}7$. Then, we give the full details on solving the Bondi accretion of $\phi^{4}$-SIDM and take the critical regime to obtain the results used in the main text. We end with two remarks: (i) dark halos as TOV configurations of SIDM; (ii) critical Bondi accretion of $\phi^{n}$-SIDM.
@@ I Forming super-Eddington SMBHs from Bondi accretion?
    Moreover, considering the accretion of cold dark matter (CDM), the local density of CDM in the inner region is estimated to be far smaller than the baryonic matter in the halos with NFW structures (Blumenthal et al., 1986; Navarro et al., 1996; Mo et al., 2010; Schaye et al., 2015). This leads to a suppressed accretion rate compared to Bondi’s prediction.
@@ II.2 Bondi accretion rate of $\phi^{4}$-SIDM
$\Big({p+\rho\over\rho_{0}}\Big)\Big(1+u^{2}-{2M\over r}\Big)=\Big({p_{\infty}+\rho_{\infty}\over\rho_{0,\infty}}\Big)$ (S18)
@@
By the consistency relations of Bondi accretion formalism, we can express $\epsilon$ as follows:
@@
This then gives the Bondi radius,
@@
Here, we identify $\rho_{0,\infty}$ the DM density inside a typical halos, denoted by $\rho_{\rm DM}$. Then, the first relation of (S27) can be replaced by
@@
where the typical virial overdensity $\Delta_{\rm vir}\simeq 300$ from spherical-collpase simulation Bryan and Norman (1998); White and Rees (1978); Blumenthal et al. (1984), then the Bondi radius can be expressed as

### v2

@@ Acknowledgements.
In this supplementary material, we will first discuss the possibility of forming the super-Eddington supermassive black holes (SMBH), which we mean the black holes with masses $\geq 10^{9}M_{\odot}$ formed around the redshift $z\simeq 6\textrm{--}7$. Then, we give the full details on solving the Bondi accretion of $\phi^{4}$-SIDM and take the critical regime to obtain the results used in the main text. We end with two remarks: (i) dark halos as TOV configurations of SIDM; (ii) critical Bondi accretion of $\phi^{n}$-SIDM. Besides, we also address the Goldilocks problem of $\phi^{4}$-SIDM and the possible resolutions.
@@ I Forming super-Eddington SMBHs from Bondi accretion?
    Moreover, considering the accretion of cold dark matter (CDM), the local density of CDM in the inner region is estimated to be far smaller than the baryonic matter in the halos with NFW structures (Blumenthal et al., 1986; Navarro et al., 1996; Mo et al., 2010; Schaye and others, 2015). This leads to a suppressed accretion rate compared to Bondi’s prediction.
@@ II.2 Bondi accretion rate of $\phi^{4}$-SIDM
$\Big({p+\rho\over\rho_{0}}\Big)^{2}\Big(1+u^{2}-{2M\over r}\Big)=\Big({p_{\infty}+\rho_{\infty}\over\rho_{0,\infty}}\Big)^{2}$ (S18)
@@
By the consistency relations of Bondi accretion formalism, we can express $\epsilon$ as follows [^7] The steps are as follows. First, combining (S12) and (S9), we obtain $\rho_{0,s}={2(1+3a_{s}^{2})^{3/2}\rho_{0,\infty}\over 9a_{s}^{2}-1}$. We further expand the above by using (S24) to get $\rho_{0,s}={16\rho_{0,\infty}\over 27\sqrt{3}\epsilon}+{\cal O}(\epsilon^{0})$, thus second equality of (S27). Next, using (S9) to solve $\rho_{\infty}$ in terms of $a_{\infty}$, then expand it in order of $\epsilon$ by using (S21) and (S24), we obtain $\rho_{\infty}={9\rho_{B}\epsilon\over 2}+{\cal O}(\epsilon^{2})$, thus the first equality of (S27). Finally, expanding $\rho_{\infty}$ of (S21) by using (S24) to obtain $a^{2}_{\infty}={9\epsilon\over 16}+{\cal O}(\epsilon^{2})$, thus the third equality of (S27).:
@@
Plugging the relation given by the third equality of (S27) into the defining (S13), we can obtain the following expression of Bondi radius,
@@
Here, we identify $\rho_{0,\infty}$ the DM density inside a typical halos, denoted by $\rho_{\rm DM}$. Moreover, $p_{\infty}\sim{\cal O}(\epsilon^{2})$ implies $p_{\infty}\rightarrow 0$. This is a natural boundary condition as the halo is of finite size, beyond which there is no DM. Then, the first relation of (S27) can be replaced by
@@
using which the Bondi radius given in (S28) becomes
$r_{B}\simeq{81\over 16}{\rho_{B}\over\rho_{\rm DM}}M\;.$ (S31)
@@
where the typical virial overdensity $\Delta_{\rm vir}\simeq 300$ from spherical-collpase simulation Bryan and Norman (1998); White and Rees (1978); Blumenthal et al. (1984), then the Bondi radius of (S31) can be expressed as
@@ II.3 Two remarks
III Goldilocks problem and its resolution by modifying $\phi^{4}$-SIDM
As discussed at the end of the Introduction, the velocity-independent $\sigma\over m$} of $\phi^{4}$-SIDM cannot help resolve the Goldilocks problem: galaxy-scale constraints require higher values of $\sigma\over m$}, but the cluster scale constraints require lower values. Instead, the Goldilocks issue can be resolved by considering SIDM with velocity-dependent cross-sections, which are large at low velocities (dwarf galaxies) and small at high velocities (clusters).
One can achieve the required cross-sections by appropriately appending additional ingredients to our SIDM model without disrupting rapid Bondi accretion in the critical regime. The simplest way is to introduce another SIDMs with the required velocity cross-sections to yield a Goldilocks zone. Unlike our $\phi^{4}$-SIDM, these new SIDMs will have negligible contributions to the Bondi accretion rate in comparison to the critical $\phi^{4}$-SIDM.
An alternative option to adding new SIDMs is to introduce a mediator scalar field $\varphi$ with the cubic interaction
${\cal L}^{\prime}={1\over 2}(\partial\varphi)^{2}-m_{\varphi}^{2}\varphi^{2}-\lambda_{3}\varphi\phi^{2}$ (S51)
which can be shown to yield a Goldilocks zone by appropriate tuning $\lambda_{3}$ and $m_{\varphi}$, e.g., by some resonant self-interaction as indicated in Tsai et al. (2022). Assuming $m_{\varphi}\gg m$, we can integrate out $\varphi$ to obtain an effecrtive $\phi^{4}$ coupling. This then change $\lambda$ to
$\lambda^{\prime}(p)=\lambda+{\cal O}(1){\lambda_{3}^{2}\over p^{2}+m^{2}_{\varphi}}\;.$ (S52)
For simplicity, we have written the effective coupling in the momentum space with $p$ denoting the off-shell 4-momentum of $\varphi$, or 4-momentum exchange of $\phi$ scattering. This is an effective velocity-dependent quartic interaction that can thus be used to resolve the Goldilocks issue, i.e., larger/smaller $\lambda^{\prime}(p)$ (cross-sections) for smaller/larger $p^{2}$ (velocities). On the other hand, when deriving the EoS of the $\phi$-field perfect fluid under self-gravity by taking the homogenous limit ($p\rightarrow 0$) in the gradient expansion, e.g., see Appendix A of Zhang et al. (2023), we will arrive at the same EoS but replacing $\lambda$ by $\lambda^{\prime}(p=0)$. However, this replacement will not affect our discussion of the critical Bondi accretion of $\phi^{4}$-SIDM by appropriately tuning $\lambda_{3}$ and $m_{\varphi}$.